OpenStudy (anonymous):

lim x->1 (sin(x-1))/(x^2+x-2) I factored the denominator. Now what?

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

(x+2)(x-1)

OpenStudy (anonymous):

shhhh. You didn't see that.

OpenStudy (anonymous):

so you got \[\lim_{x\to 1}\frac{\sin(x-1)}{x-1}\times \frac{1}{x+2}\] first part you should remember and the second is obvious

OpenStudy (anonymous):

but it's lim x->1

OpenStudy (anonymous):

I thought the identity was only x->0

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

it is clear to you that as \(x\to 1\) you have \(x-1\to 0\) right?

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

in other words, \[\lim_{x\to 0}\frac{\sin(x)}{x}=\lim_{x\to 1}\frac{\sin(x-1)}{x-1}\] it is just a change of variable

OpenStudy (anonymous):

wat. witchcraft.

OpenStudy (anonymous):

Okay, then. 1/3

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

witchcraft for sure, or saying "put \(x-1=y\) " either way yw